# Broken symmetry solutions in one-dimensional lattice models via many-body perturbation theory

### v1

@@ III.2 Self-consistency
Since the Dyson equation is solved iteratively, the accuracy of the algorithm is crucial to obtain faithful solutions of the self-consistent cycle. Considering a non-interacting $G_{0}$ starting point, at each iteration the self-energy is evaluated in terms of its SOP representation and the Dyson equation is solved via the algorithmic inversion method described in Sec. III. The self-consistency loop is stopped whenever each matrix element $\gamma_{ij}$ of the one-body density matrix changes less than a given threshold $\epsilon$ between two successive iterations:
@@
Typical thresholds are within $10^{-8}<\epsilon<10^{-6}$.
In order to keep the number of poles under control we adopt a decimation procedure consisting in merging and pruning the poles of the dynamical operators, which is commonly employed when working with the meromorphic representation of dynamical operators Puig von Friesen et al. (2010); Di Sabatino et al. (2021). If the distance between two poles is smaller than a given threshold $\delta_{M}$, then the poles are merged into a new pole, whose residue is given by the sum of the residues of the former poles and whose position is given by a weighted average of the previous poles, using weights $w_{s}$ related to the trace of the residues by
@@
where the $A^{s}$ are the residues of the dynamical operator as in Eq. (8). If all the matrix elements of the residue are instead smaller than the pruning threshold $\delta_{P}$ the pole is then eliminated and its residue is redistributed among the two nearest non-pruned poles with weights $w_{s}$ given again by Eq. (15). Since the pruning and merging operations may alter the value of the ranks of the residues, the $\text{Tr}_{\omega}\text{Ln}$ term appearing in Eq. (III.1) is always evaluated before these operations are performed.
In order to achieve self-consistency, it is generally necessary to mix dynamical operators calculated at different iterations of the self-consistency loop. A commonly adopted strategy is that of linearly mixing the GFs between two successive iterations Puig von Friesen et al. (2010); Joost et al. (2022); Honet et al. (2022). Alternatively, we also adopt a linear mixing between the self-energies coming from two successive iterations Koval et al. (2014):
@@
where the mixing parameter is $10<\alpha<$ 1 and the mixed GF is then evaluated by solving a Dyson equation with the mixed self-energy.
The Sham-Schlüter equation, Eq. (II.2), is solved in practice after the solution of the self-consistent Dyson equation. The SSE on a discrete lattice defines a linear system of equations having $v_{KS}(\mathbf{r})$ as unknown. This is solved self-consistently by starting from a guess of the KS potential (e.g. the local entries of the static part of the electron-electron self-energy) and solving the system by iterations until self-consistency in the density matrix is obtained. At each iteration of the SSE loop the highest “occupied” pole of $G_{KS}$ is aligned to that of the corresponding $G$ to ensure that they share the same Fermi level.
@@ IV Validation
We start by comparing the results of our work against the available literature. In particular, the Hubbard model constitutes an optimal playground to study the effects of electronic correlation, where self-consistent calculations have already been performed. The Hubbard model on a lattice Hubbard (1963); Gutzwiller (1963); Kanamori (1963) includes only a local electron-electron repulsion beyond the independent-particle hopping and its Hamiltonian is given by
@@
In this work we focus on the one-dimensional Hubbard model with only one orbital per site and at half-filling. We compare our results to those in Refs. Puig von Friesen et al. (2010); Di Sabatino et al. (2021); Joost et al. (2022). To this aim we stress that, in the case of small Hubbard chains, an exact treatment is also numerically feasible through exact diagonalization, which is reported in the Supplemental Material sup . An analytical treatment of the system also shows that no metal-insulator transition (MIT) is found in the case of chains with an even number of sites and at half-filling Lieb and Wu (1968, 2003b).
@@ IV.1 Hubbard dimer: nonmagnetic solution
Starting from the half-filled Hubbard dimer we compare the spectral features of quasiparticle (QP) poles and satellites against the results of Di Sabatino et al. Di Sabatino et al. (2021) in the case of a self-consistent GW calculation. The $Z$ factor for a given pole of the GF is $Z_{s}=\langle f_{s}|f_{s}\rangle=\text{Tr}[A^{s}]$, with $A^{s}$ defined according to Eq. (8). The exact one-particle GF of the symmetric Hubbard dimer at half filling has two occupied and two empty poles (one quasi-particle and one satellite in each case), symmetric with respect to the Fermi level (here set at zero), because of particle-hole symmetry. When comparing with the numerically computed GFs (which are in general approximate because of the chosen self-energy), we consider as QP the pole having the largest $Z$ factor, while we take as satellite the pole having the second largest $Z$ factor.
@@
This work Ref. Di Sabatino et al. (2021) This work Ref. Di Sabatino et al. (2021)
@@
This work Ref. Di Sabatino et al. (2021) This work Ref. Di Sabatino et al. (2021)
@@
Table 1: Hubbard dimer treated at the GW level compared with Ref. Di Sabatino et al. (2021). We report the QP (upper panel) and satellite (lower panel) pole positions and their $Z$ factors. The Fermi level is shifted at $E_{F}=0$. The thresholds used in the calculations are $\delta_{M}=10^{-2}$ and $\delta_{P}=10^{-6}$, with a convergence threshold of $\epsilon=10^{-8}$. Frequencies are reported in units of $t$. We have highlighted in bold text the digits for which our results agree with those of the reference work.
Overall, we obtain an excellent numerical agreement with the work of Ref. Di Sabatino et al. (2021), where a meromorphic representation of the dynamical operators is also employed, with the poles of the GF found via a root finding algorithm. The agreement in the case of the $Z$ factors is almost perfect, while energy differences around $10^{-3}$ (in units of $t$) are found for the QP poles and energy differences of $10^{-2}$ are found for the satellites in the regime of large $U/t$.
@@ IV.2 Hubbard dimer: antiferromagnetic solution
Within this Section we discuss in detail the features of the broken-symmetry solutions obtained in the Hubbard dimer. In passing we stress that the exact ground state of a one-dimensional Hubbard chain at half-filling and with an even number of sites is always NM and with a homogeneous charge density, as can be deduced from exact diagonalization results Joost et al. (2022); Honet et al. (2022). AFM solutions for the ground state are then unphysical, in the sense that they violate the symmetry of the exact solution. Nevertheless the study of these solutions is still relevant, as it allows one to obtain better estimates for thermodynamic and spectral quantities, such as the total energy and energy gap Joost et al. (2022); Honet et al. (2022).
Using our fully self-consistent implementation we investigate the symmetry breaking, specializing to the case of GW. In order to quantitatively characterize AFM solutions, and following Ref. Honet et al. (2022), we introduce the following long-range order parameter:
@@
We observe that in the case of the Hubbard dimer HF gives an AFM ground state for values of $U/t>2$, similarly to what found in Ref. Joost et al. (2022). We also stress that, within HF, no AFM self-consistent solution is actually found for $U/t<2$, while for $U/t>2$ we can find either NM or AFM solutions by changing the symmetry of the starting GF of the self-consistency loop.
@@
The GW calculations are initialized from the HF self-consistent solutions obtained at the corresponding value of $U/t$. We observe that, at variance with Ref. Honet et al. (2022), even if the self-consistency loop starts from an AFM GF, it does not necessarily converge to an AFM solution, while a NM starting GF always restricts the symmetry of the ground state. These results are reported in Fig. 1, where we study the energy difference (left axis) between the unrestricted symmetry (US) calculation with respect to the NM one. In Fig. 1 we also report the order parameter (right axis) of the self-consistent GF at the end of the loop for the US case. In the case of GW we find that an AFM ground state is found for $U/t>4.21$, while there is no AFM solution below that value, signaled by the vanishing order parameter $M^{LRO}$ in Fig. 1. We notice that the AIM-SOP implementation is capable of high accuracy in resolving total energy differences also below $10^{-3}$ (units of $t$ are used throughout). Furthermore the presence of an AFM solution at $U/t\simeq 4.21$ is signaled by the simultaneous onset of the order parameter $M^{LRO}$ at the same value of $U/t$ where the energy difference $E-E_{NM}$ becomes negative.
@@
In order to assess the stability of our AIM-SOP implementation, we study how the total energy and the order parameter change within the self-consistency loop. In Fig. 2 we report the values of the total energy and of the order parameter for the half-filled Hubbard dimer obtained in GW for an US calculation. We observe that when the stopping condition of the loop is reached, the Klein functional of Eq. (4) agrees with the energy value returned from the Galitskii-Migdal formula, as expected. In general the agreement between different functionals (or with the GM formula) is a non-trivial result and it can also be taken as a diagnostic that self-consistency has been achieved Puig von Friesen et al. (2010).
We note that for $U/t<4.21$ the loop converges towards a NM solution, by slowly reducing the order parameter at each iteration. Instead, above the value of $U/t>4.21$ the loop converges to an AFM self-consistent solution. The order parameter is finite in this situation, but it is smaller than that of the corresponding HF AFM solution. Another relevant consideration in this case is that the total energy reaches a stable value earlier with respect to the order parameter, so that relative differences in the order parameter for this system may be more evident than differences in total energies. Overall, the total energy and the order parameter evolve smoothly during the self-consistency loop, assuring the stable behaviour of our AIM-SOP implementation. At variance to GW, when using a 2B self-energy AFM solutions can be stabilized even when they are not the ground state of the system. Broken symmetry solutions of the Hubbard dimer found using the 2B are discussed in detail in the Supplemental Material sup , where we also comment on the difference with respect to GW.
We stress that within our AIM-SOP implementation we only manage to find at most one self-consistent solution for each symmetry (NM or AFM) at any given value of $U/t$. Thus multiple solutions with the same symmetry have not been found. Furthermore, we remark that the poles of the GF at each step of the self-consistency loop are always found on the real axis if the starting GF has poles on the real axis, and the residues of the GF are always positive semi-definite (PSD) for the self-energy approximations here considered. This implies that at each step of the loop the GF is compatible with a Lehmann representation Lehmann (1954), and that all the expected analytical properties are satisfied Ferretti et al. (2024). In particular, as analyzed in Ref. Ferretti et al. (2024), if the input self-energy is well-behaved, i.e. if it has Hermitian PSD residues, real poles, and Hermitian asymptotic value $\Sigma_{0}$, the AIM-SOP procedure leads to a Hermitian effective AIM Hamiltonian, out of which the resulting GF is also well-behaved (PSD residues and real poles). The reverse is also true Ferretti et al. (2024).
@@ IV.3 Hubbard chains: spectral properties
We now move to study the spectral features of longer Hubbard chains, where the QP and satellite structure is richer with respect to that of the Hubbard dimer. The results of our AIM-SOP implementation are compared with those of Refs. Puig von Friesen et al. (2010); Joost et al. (2022). We start by considering a $N=6$ Hubbard chain at half-filling, with $U/t=4$ and open boundary conditions, and focus on the NM solution.
@@
The spectral density projected on the first site of the chain is reported in Fig. 3. We show the spectral density obtained with the different approximations and also the exact result obtained via exact diagonalization. Overall we obtain an excellent agreement with the work of von Friesen et al. Puig von Friesen et al. (2010) for the HF, 2B, and GW levels of theory. In particular, HF gives the same spectral density of the non-interacting system, consisting of three main occupied peaks (plus particle-hole symmetry). Besides this energy region, the 2B self-energy gives satellites with larger spectral weight, while in GW a structureless spectrum of satellites appears.
As also stressed in Ref. Puig von Friesen et al. (2010), we notice that the self-consistency of GW in longer chains and with large $U/t$ becomes slow and difficult, since this self-energy approximation gives rise to several satellites with small residues. In particular, for GW the position of the furthest satellite from the Fermi level (around $U/t=7.5$) changes weakly with the pruning threshold. The residue of this satellite is also larger with respect to the nearby ones since, due to our pruning scheme, it is receiving the spectral weight of the pruned poles far away from the Fermi level.
We then consider the case of a $N=8$ Hubbard chain with periodic boundary conditions and $U/t=4$. The same system has been considered by Joost et al. in Ref. Joost et al. (2022), where a real-time implementation has been employed. We focus here on an AFM solution found with the self-consistent 2B self-energy. The density of states (DOS) of this chain is reported in Fig. 4.
@@
We find an excellent agreement with respect to the DOS of Ref. Joost et al. (2022), while we also observe a very good agreement for the value of the total energy: in our case we obtain $E^{LWK}_{tot}/t=-4.069(0)$ and $E^{GM}_{tot}/t=-4.069(3)$, against $E^{ref}_{tot}/t=-4.08$. We notice how the AIM-SOP implementation is also able to resolve well the presence of the satellites for the 2B approximation between $\omega=-6.0t$ and $\omega=-4.0t$. In passing, we note that we could not exactly reproduce the results of Ref. Honet et al. (2022). We think that a different self-consistency condition between our implementation and that of Ref. Honet et al. (2022) is among the possible causes of discrepancy. A detailed study of a $N=6$ Hubbard chain, both in the NM and unrestricted-symmetry calculations, is reported in the Supplemental Material sup .

### v2

@@ III.2 Self-consistency
Since the Dyson equation is solved iteratively, the accuracy of the algorithm is crucial to obtain faithful solutions of the self-consistent cycle. Considering a non-interacting $G_{0}$ starting point, at each iteration the self-energy is evaluated in terms of its SOP representation and the Dyson equation is solved via the algorithmic-inversion method described in Sec. III. The self-consistency loop is stopped whenever each matrix element $\gamma_{ij}$ of the one-body density matrix changes less than a given threshold $\epsilon$ between two successive iterations:
@@
Typical thresholds are $10^{-8}<\epsilon<10^{-6}$.
In order to keep the number of poles under control we adopt a decimation procedure consisting in merging and pruning the poles of the propagators or self-energies, which is commonly employed when working with the meromorphic representation of dynamical operators [28, 29]. If the distance between two poles is smaller than a given threshold $\delta_{M}$, the poles are merged into a new pole, with the residue given by the sum of the residues of the former poles, and located at a weighted average of the previous poles. The weights $w_{s}$ are related to the trace of the residues by
@@
where the $A^{s}$ are the residues of the dynamical operator as in Eq. (8). If all the matrix elements of the residue are instead smaller than the pruning threshold $\delta_{P}$, the pole is then eliminated and its residue is redistributed among the two nearest non-pruned poles, with weights $w_{s}$ given again by Eq. (15). Since the pruning and merging operations may alter the value of the ranks of the residues, the $\text{Tr}_{\omega}\text{Ln}$ term appearing in Eq. (III.1) is always evaluated before these operations are performed.
In order to achieve self-consistency, it is generally necessary to mix dynamical operators calculated at different iterations of the self-consistency loop. A commonly adopted strategy is that of linearly mixing the GFs between two successive iterations [28, 7, 8]. Alternatively, we also adopt a linear mixing between the self-energies coming from two successive iterations [63]:
@@
where the mixing parameter is $10<\alpha<$ 1 and the mixed GF is then evaluated by solving a Dyson equation with the mixed self-energy. According to our simulations, the $G$-mixing scheme tends to converge faster than $\Sigma$-mixing, at least for the systems studied in this work.
The Sham-Schlüter equation, Eq. (II.2), is solved in practice after the solution of the self-consistent Dyson equation. The SSE on a discrete lattice defines a linear system of equations having $v^{\text{KS}}(\mathbf{r})$ as unknown. This is solved self-consistently by starting from a guess of the KS potential (e.g., the local entries of the static part of the electron-electron self-energy) and solving the system by iterations until self-consistency in the density matrix is obtained. At each iteration of the SSE loop the highest “occupied” pole of $G^{\text{KS}}$ is aligned to that of the corresponding $G$ to ensure that they share the same Fermi level.
@@ IV Validation
We start by comparing the results of our work against the available literature. In particular, the Hubbard model constitutes an optimal playground to study the effects of electronic correlation, where self-consistent MBPT calculations have already been performed. The Hubbard model on a lattice [6, 64, 65] includes only a local electron-electron repulsion beyond the independent-particle hopping and its Hamiltonian is given by
@@
In this work we focus on the one-dimensional Hubbard model with only one orbital per site and at half-filling. We compare our results to those in Refs. [28, 29, 7]. To this aim we stress that, in the case of small Hubbard chains, an exact treatment is also numerically feasible through exact diagonalization, which is reported in the Supplemental Material [42]. An exact analytical treatment of the system also shows that no metal-insulator transition (MIT) is found in the case of chains with an even number of sites and at half-filling [12, 13].
@@ IV.1 Hubbard dimer: nonmagnetic solution
Starting from the half-filled Hubbard dimer, we compare the spectral features of quasiparticle (QP) poles and satellites against the results of Di Sabatino et al. [29] in the case of a self-consistent GW calculation. The $Z$ factor for a given pole of the GF is $Z_{s}=\langle f_{s}|f_{s}\rangle=\text{Tr}[A^{s}]$, with $A^{s}$ defined according to Eq. (8). The exact one-particle GF of the symmetric Hubbard dimer at half filling has two occupied and two empty poles (one quasi-particle and one satellite in each case), symmetric with respect to the Fermi level (here set at zero), because of particle-hole symmetry. When comparing with the numerically computed GFs (which are in general approximate because of the chosen self-energy), we consider as QP the pole having the largest $Z$ factor, while we take as satellite the pole having the second largest $Z$ factor.
@@
This work Ref. [29] This work Ref. [29]
@@
This work Ref. [29] This work Ref. [29]
@@
Table 1: Hubbard dimer treated at the GW level compared with Ref. [29]. We report the QP (upper panel) and satellite (lower panel) pole positions and their $Z$ factors. The Fermi level is shifted at $E_{F}=0$. The thresholds used in the calculations are $\delta_{M}=10^{-2}$ and $\delta_{P}=10^{-6}$, with a convergence threshold of $\epsilon=10^{-8}$. Frequencies are reported in units of $t$. We have highlighted in bold text the digits for which our results agree with those of the reference work.
Overall, we obtain an excellent numerical agreement with the work of Ref. [29], where a meromorphic representation of the dynamical operators is also employed, with the poles of the GF found via a root finding algorithm. The agreement in the case of the $Z$ factors is almost perfect, while energy differences around $10^{-3}$ (in units of $t$) are found for the QP poles and energy differences of $10^{-2}$ are found for the satellites in the regime of large $U/t$.
@@ IV.2 Hubbard dimer: antiferromagnetic solution
Within this Section we discuss in detail the features of the broken-symmetry solutions obtained in the Hubbard dimer. In passing we stress that the exact ground state of a one-dimensional Hubbard chain at half-filling and with an even number of sites is always NM and with a homogeneous charge density [12, 13] , as also found by high-accuracy and exact diagonalization methods [10, 66, 7, 8]. AFM solutions for the ground state are then unphysical, in the sense that they violate the symmetry of the exact solution. Nevertheless, the study of these solutions is still relevant, as it may allow one to obtain better estimates for thermodynamic and spectral quantities, such as the total energy and energy gap. In fact, following the observation of Löwdin [67], in the case of Hartree-Fock calculations lifting the symmetry constraints on the eigenfunctions may improve the variational flexibility of the method. The same consideration has been extended to the case of dynamical self-energies such as 2B [7] and GW [8].
Using our fully self-consistent implementation we investigate the symmetry breaking, specializing to the case of GW. In order to quantitatively characterize AFM solutions, and following Ref. [8], we introduce the following long-range order parameter:
@@
We observe that in the case of the Hubbard dimer HF gives an AFM ground state for values of $U/t>2$, similarly to what found in Ref. [7]. We also stress that, within HF, no AFM self-consistent solution is actually found for $U/t<2$, while for $U/t>2$ we can find either NM or AFM solutions by changing the symmetry of the starting GF of the self-consistency loop.
@@
The GW calculations are initialized from the HF self-consistent solutions obtained at the corresponding value of $U/t$. We observe that, at variance with Ref. [8], even if the self-consistency loop starts from an AFM GF, it does not necessarily converge to an AFM solution, while a NM starting GF always restricts the symmetry of the ground state. These results are reported in Fig. 1, where we study the energy difference (left axis) between the unrestricted symmetry (US) calculation with respect to the NM one. In Fig. 1 we also report the order parameter (right axis) of the self-consistent GF at the end of the loop for the US case. In the case of GW we find that an AFM ground state is found for $U/t>4.21$, while there is no AFM solution below that value, signaled by the vanishing order parameter $M^{LRO}$ in Fig. 1. We notice that the AIM-SOP implementation is capable of high accuracy in resolving total energy differences also below $10^{-3}$ (units of $t$ are used throughout). Furthermore the presence of an AFM solution at $U/t\simeq 4.21$ is signaled by the simultaneous onset of the order parameter $M^{LRO}$ at the same value of $U/t$ where the energy difference $E-E_{NM}$ becomes negative. The AFM solution persists also in the atomic limit, where it gives rise to a finite energy gap that scales with $U$, consistently with the exact result and at variance with the metallic NM solution within GW [12, 68, 69].
@@
In order to assess the stability of our AIM-SOP implementation, we study how the total energy and the order parameter change within the self-consistency loop. In Fig. 2 we report the values of the total energy and of the order parameter for the half-filled Hubbard dimer obtained in GW for an US calculation. We observe that when the stopping condition of the loop is reached, the Klein functional of Eq. (4) agrees with the energy value returned from the Galitskii-Migdal formula, as expected. In general the agreement between different functionals (or with the GM formula) is a non-trivial result and it can also be taken as a diagnostic that self-consistency has been achieved [28].
We note that for $U/t<4.21$ the loop converges towards a NM solution, by slowly reducing the order parameter at each iteration. Instead, above the value of $U/t>4.21$ the loop converges to an AFM self-consistent solution. The order parameter is finite in this situation, but it is smaller than that of the corresponding HF AFM solution. Another relevant consideration in this case is that the total energy reaches a stable value earlier with respect to the order parameter, so that relative differences in the order parameter for this system may be more evident than differences in total energies. Overall, the total energy and the order parameter evolve smoothly during the self-consistency loop, assuring the stable behaviour of our AIM-SOP implementation. At variance to GW, when using a 2B self-energy AFM solutions can be stabilized even when they are not the ground state of the system. Broken symmetry solutions of the Hubbard dimer found using the 2B are discussed in detail in the Supplemental Material [42], where we also comment on the difference with respect to GW.
We stress that within our AIM-SOP implementation, whenever starting from a non-interacting $G_{0}$ having real poles and positive semi-definite (PSD) residues, we only manage to find at most one self-consistent solution for each symmetry (NM or AFM) at any given value of $U/t$. Thus multiple solutions with the same symmetry have not been found, neither considering different starting points as discussed for the 2B self-energy in the Supplemental Material [42], nor adopting different mixing schemes as discussed in Sec. III.2. Furthermore, we remark that the poles of the GF at each step of the self-consistency loop are always found on the real axis if the starting GF has poles on the real axis, and the residues of the GF are always PSD for the self-energy approximations here considered. This implies that at each step of the loop the GF is compatible with a Lehmann representation [60], and that all the expected analytical properties are satisfied [40]. In particular, as analyzed in Ref. [40], if the input self-energy is well-behaved, i.e. if it has Hermitian PSD residues, real poles, and Hermitian asymptotic value $\Sigma_{0}$, the AIM-SOP procedure leads to a Hermitian effective AIM Hamiltonian, out of which the resulting GF is also well-behaved (PSD residues and real poles). The reverse is also true [40].
@@ IV.3 Hubbard chains: spectral properties
We now move to study the spectral features of longer Hubbard chains, where the QP and satellite structure is richer with respect to that of the Hubbard dimer. The results of our AIM-SOP implementation are compared with those of Refs. [28, 7]. We start by considering a $N=6$ Hubbard chain at half-filling, with $U/t=4$ and open boundary conditions, and focus on the NM solution.
@@
The spectral density projected on the first site of the chain is reported in Fig. 3. We show the spectral density obtained with the different approximations and also the exact result obtained via exact diagonalization. Overall we obtain an excellent agreement with the work of von Friesen et al. [28] for the HF, 2B, and GW levels of theory. In particular, HF gives the same spectral density of the non-interacting system, consisting of three main occupied peaks (plus particle-hole symmetry). Besides this energy region, the 2B self-energy gives satellites with larger spectral weight, while in GW a structureless spectrum of satellites appears.
As also stressed in Ref. [28], we notice that the self-consistency of GW in longer chains and with large $U/t$ becomes slow and difficult, since this self-energy approximation gives rise to several satellites with small residues. In particular, for GW the position of the furthest satellite from the Fermi level (around $\omega/t=7.5$) changes weakly with the pruning threshold. The residue of this satellite is also larger with respect to the nearby ones since, due to our pruning scheme, it is receiving the spectral weight of the pruned poles far away from the Fermi level.
We then consider the case of a $N=8$ Hubbard chain with periodic boundary conditions and $U/t=4$. The same system has been considered by Joost et al. in Ref. [7], where a real-time implementation has been employed. We focus here on an AFM solution found with the self-consistent 2B self-energy. The density of states (DOS) of this chain is reported in Fig. 4.
@@
We find an excellent agreement with respect to the DOS of Ref. [7], while we also observe a very good agreement for the value of the total energy: in our case we obtain $E^{LWK}_{\text{tot}}/t=-4.069(0)$ and $E^{GM}_{\text{tot}}/t=-4.069(3)$, against $E^{\text{ref}}_{\text{tot}}/t=-4.08$. We notice how the AIM-SOP implementation is also able to resolve well the presence of the satellites for the 2B approximation between $\omega=-6.0t$ and $\omega=-4.0t$. In passing, we note that we could not exactly reproduce the results of Ref. [8]. We believe that a different self-consistency condition may be among the possible causes of discrepancy. A detailed study of a $N=6$ Hubbard chain, both in the NM and unrestricted-symmetry calculations, is reported in the Supplemental Material [42].